OpenStudy (anonymous):

can someone please help me with factoring! MEDAL AWARDED 22x^2+50x+12

OpenStudy (anonymous):

How would we begin o factor this out?

OpenStudy (anonymous):

using the x-box method i believe

OpenStudy (anonymous):

\[22x^2+50x+12\]

OpenStudy (anonymous):

it's 6*44

OpenStudy (anonymous):

2(11x+3)(x+2)

OpenStudy (anonymous):

6+44=50, and 6*44=264

OpenStudy (anonymous):

There is not 264 in your original un-factored polynomial

OpenStudy (anonymous):

\[\large 2(11x^2+25x+6)\] First step, factor out 2. Now find factors of 11 x 6 whose sum is 25. Factoring the quadratic \[\large 11x^2+25x+6 ~ => ~ 11x^2+22x+3x+6 \] Because the product of 11 and 6 is 66. The factors of 66 which sum to 25 are 3 and 22. So , \[\large 11x^2+22x+3x+6 ~ => ~ 2(11x+3)+x(11x+3)\] Factor out the common terms now \[\large 2(11x+3)(x+2)\]

OpenStudy (anonymous):

how would to factor the 2(11x+3)(x+2)?

OpenStudy (anonymous):

You don't that's the final answer.

OpenStudy (anonymous):

It can not be "factored", the factoring shows the polynomial function of higher degrees to lower as you may notice there is no longer an x^2

OpenStudy (anonymous):

but don't you need to find the roots? the x values setting 2(11x+3)(x+2)=0?

OpenStudy (anonymous):

Is it asking you to find the roots?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright so basically you just solve for x then, 11x+3 = 0, and x+2 = 0, doing that you get x = -2, and x = -3/11 :)!

OpenStudy (anonymous):

:) thanks so much!

OpenStudy (anonymous):

Np ^.^